OpenStudy (anonymous):

Use the T-table below to determine 3 solutions to this equation: 4x - 2y = 8 all the X's are 0 1 -2

OpenStudy (anonymous):

I have absolutely no idea how to do T-Tables D:

OpenStudy (anonymous):

given eqn is 4x -2y = 8 put x = 0 in it and u get 4(0) - 2y =8 -2y = 8 y = 8/-2 y = -4 so first solution is (0, -4)

OpenStudy (anonymous):

given eqn is 4x -2y = 8 put x = 1 in it and u get 4(1) - 2y =8 4 -2y = 8 -2y = 8 -4 -2y = 4 y = 4/-2 y = -2 so second solution is (1, -2)

OpenStudy (anonymous):

similarly given eqn is 4x -2y = 8 put x = -2 in it and u get 4(-2) - 2y =8 -8 -2y = 8 -2y = 8 + 8 -2y = 16 y = 16/-2 y = -8 so third solution is (-2, -8)